OpenStudy (anonymous):

Find the equations for the lines through the point (a, c) that are parallel to and perpendicular to the line y = mx + b. Assume m ≠ 0. Parallel y= perpendicular y=

OpenStudy (anonymous):

y-c=m(x-a) y=mx-am+c eq===> paralel

OpenStudy (anonymous):

perpendicular y-c=-1/m(x-a) my-mc=-x+a my=-x+mc+a y=1/m(-x+mc+a)

OpenStudy (anonymous):

can you please explain how to do the perpendicular one?

OpenStudy (anonymous):

for perpendicular one the slope will be -1/m

OpenStudy (anonymous):

rest will be same

OpenStudy (anonymous):

why is it -1/m ???

OpenStudy (anonymous):

because the product of slope of perpendicular lines is equal to negative one